OpenStudy (anonymous):

Write and solve the proportion: 3 is to 12 as w is to 8.

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow 3:12::w:8 }\) \(\Large \color{purple}{\rightarrow {3 \over 12} = {w \over 8} }\) \(\Large \color{purple}{\rightarrow 8 * 3 = 12w }\) \(\Large \color{purple}{\rightarrow 24 = 12w }\) Solve for w now.

OpenStudy (anonymous):

\[\frac{3}{12}=\frac{w}{8} \] \[ \frac{1}{4}=\frac{w}{8} \] \[\frac{8}{4}=w \] \[w=2 \]

OpenStudy (anonymous):

thankssss

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

suck ma left nutt pussy